# IRIF, Université Paris Cité, France

### v1

@@ Abstract
Neural network model extraction has emerged in recent years as an important security concern, as adversaries attempt to recover a network’s parameters via black-box queries. A key step in this process is signature extraction, which aims to recover the absolute values of the network’s weights layer by layer. Prior work, notably by Carlini et al. (2020), introduced a technique inspired by differential cryptanalysis to extract neural network parameters. However, their method suffers from several limitations that restrict its applicability to networks with a few layers only. Later works focused on improving sign extraction, but largely relied on the assumption that signature extraction itself was feasible.
In this work, we revisit and refine the signature extraction process by systematically identifying and addressing for the first time critical limitations of Carlini et al.’s signature extraction method. These limitations include rank deficiency and noise propagation from deeper layers. To overcome these challenges, we propose efficient algorithmic solutions for each of the identified issues, greatly improving the efficiency of signature extraction. Our approach permits the extraction of much deeper networks than was previously possible. We validate our method through extensive experiments on ReLU-based neural networks, demonstrating significant improvements in extraction depth and accuracy. For instance, our extracted network matches the target network on at least 95% of the input space for each of the eight layers of a neural network trained on the CIFAR-10 dataset, while previous works could barely extract the first three layers. Our results represent a crucial step toward practical attacks on larger and more complex neural network architectures.
Keywords:
ReLU-based neural networks signature extraction weight-recovery
@@ 1 Introduction
Neural Networks (NNs) are a class of machine learning models composed of layers of interconnected units (or neurons) that transform input data and learn to recognize patterns through a training process.
@@
Deep Neural Networks (DNNs) are a subset of NNs with multiple hidden layers, enabling them to model highly complex functions. Their depth allows them to capture complex patterns and abstract features, making them particularly effective for large-scale and high-dimensional data. DNNs have become indispensable in various applications, including computer vision (e.g., image recognition, and video analysis), natural language processing, automated medical diagnostics, and fraud detection, among others.
However, training DNNs requires large datasets, significant computational resources, and carefully fine-tuned algorithms [5, 27]. As a result, trained DNNs have become valuable intellectual assets, making them attractive targets for attackers seeking to extract the model rather than invest in training their own.
@@
In many cases, DNNs are deployed as cloud-based or online services, allowing users to interact with them without direct access to their internal parameters [25, 1]. A natural question, therefore, is how well different DNNs resist attacks in this black-box setting, where an adversary can query the network using random or carefully chosen inputs—potentially adaptively—and exploit the observed outputs to reconstruct either the exact internal parameters or an approximation sufficient to construct a functionally equivalent network.
Extracting the parameters of a neural network from its black-box implementation has been extensively studied since 2016 [26, 20, 24, 14, 8, 11, 19, 13]. The difficulty of the extraction problem varies depending on what the adversary is allowed to observe in the output. The most realistic but challenging scenario for the attacker is when only the final label is available. This label indicates the most probable category for a given input, without exposing additional confidence scores or intermediate values. This setting is notably referred to as S1 in [7]. At the other end of the spectrum, the easiest scenario, called S5 in [7], allows the attacker to access the raw output of the neural network before the normalization phase, providing much richer information for reconstruction.
@@
A significant step in model extraction research was made by Carlini et al. [8] in 2020 within the S5 scenario, specifically for networks using ReLU as the activation function, where they reframed the problem from a cryptanalytic perspective. They observed that DNNs share key similarities with block ciphers, including their iterative structure, alternating linear and non-linear layers, and the use in parallel of a small function (S-box in block ciphers, activation function in neural networks) to ensure non-linearity. Taking inspiration from cryptanalytic attacks on block ciphers, particularly differential cryptanalysis [4], Carlini et al. proposed a new two-step iterative approach for recovering a model’s internal parameters. Their method decomposes the recovery process into two distinct phases: signature extraction and sign extraction. In the first step, the absolute values of the weights in a layer are reconstructed (signature extraction), followed by a second step where the correct signs are determined (sign extraction). This process is applied iteratively, proceeding layer by layer. This approach enabled them to successfully extract neural networks, including a model with $100,000$ parameters trained on the MNIST digit recognition task, using only $2^{21.5}$ queries and less than an hour of computation. However, their method for extracting the signs had an exponential complexity in the number of neurons, limiting its feasibility to small networks with typically very few (at most 3) layers.
@@
In 2024, Canales-Martínez et al. [6] extended the work of Carlini et al. [8] by introducing several new algorithms for the sign-recovery step in ReLU-based neural networks. Their improved sign-extraction methods permitted them to extract the parameters of significantly larger networks than those considered in [8]. For instance, they reported successfully extracting the parameters of a neural network with 8 hidden layers of 256 neurons each—amounting to 1.2 million parameters—trained on the CIFAR-10 dataset [16] for image classification across 10 categories. It is important to note, however, that their experiments assumed the absolute values of the weights had already been recovered using the method of Carlini et al. [8] and were not implemented as full end-to-end attacks.
@@
More recently, Foerster et al. [13] conducted a detailed analysis of the signature and sign extraction procedures, implementing and benchmarking them on various ReLU-based neural networks. Their work revealed that, contrary to what was previously believed, the main practical bottleneck in these attacks lies in the signature extraction step introduced in [8]. Because of this, both [8] and [13] failed in practice to recover weight parameters beyond the third layer. Thus, although the authors of [6] reported successful extraction from deeper networks, this was only partially true, as the weights were supposed to have been extracted with  [8]’s method. Indeed, as Foerster et al. [13] demonstrated, the signature extraction step remains a major limitation, preventing all these methods from progressing beyond the early layers.
@@
As of today, no algorithm is known to recover efficiently a neural network’s parameters beyond the third layer [^1] The authors of [6] report having successfully extracted the 4th layer of a single network. However, in that network, this 4th layer is very contractive and in this setting, the issues identified in our work do not arise.. This highlights the need for significant improvements in the signature recovery process to enable practical attacks on deeper networks.
@@
In this work, we revisit the signature extraction algorithm of [8], which serves as the basis for the most important parameter extraction attacks against NNs [8, 6, 9, 7]. We start by carefully analyzing its limitations and highlighting the main problems that reduce its effectiveness. More specifically, our analysis reveals two major issues:
Impact of deeper layers on signature merging. Partial signature merging is a key step in signature recovery. While the impact of deeper layers on this process was previously considered negligible, we show that in deep neural networks, they introduce significant interference. This added noise often causes neurons to be misattributed to the wrong layer during merging.
Rank deficiency. In deeper layers, the system of equations collected by the attacker to recover a neuron’s weights may be rank-deficient, leading to an erroneous neuron recovery.
In a second step, we propose concrete algorithmic solutions to overcome both identified issues. We implemented and evaluated these solutions on neural networks trained on different datasets. Our results demonstrate, for the first time, that it is possible to extract nearly all weights from deep networks. In particular, while previous state-of-the-art signature extraction methods could not recover in practice layers beyond the third layer, our approach successfully extracts neurons from networks with up to eight hidden layers. We also applied our methods to even deeper networks with sixteen hidden layers and but we encountered runtime limitations that prevented full recovery.
2 Related Work
Deep neural network model extraction is an old and well-studied problem. The earliest work dates back to 1994, when Fefferman [12] proved that the output of a sigmoid network uniquely determines its architecture and weights, up to trivial equivalences. Later, in 2005, Lowd and Meek [18] introduced new algorithms for reverse engineering linear classifiers and applied these techniques to spam filtering, marking one of the first practical extraction attacks. The field saw renewed interest in 2016 with the work of Tramèr et al. [26], who demonstrated attacks on deployed machine learning models accessible through APIs that output high-precision confidence values (also referred to as scores), in addition to class labels. By exploiting these confidence scores, they successfully mounted attacks on various model types, including logistic regression and decision trees.
For non-linear models, [26] and later [23] introduced the notion of task accuracy extraction, where the goal is to build a model that performs well on the same decision task, achieving high accuracy on predictions—without necessarily reproducing the exact outputs of the original model. This is different from functionally equivalent extraction, where the objective is to replicate the original model’s outputs on all inputs, regardless of the ground truth. A related, but narrower concept is fidelity extraction, which focuses on replicating the model’s predictions over a specific input distribution. Jagielski et al. [15], who introduced this taxonomy, argued that learning-based approaches like [26, 23] are fundamentally limited when it comes to achieving functionally equivalent extraction.
Early attempts for functionally equivalent extraction relied on access to internal gradients of the model [20], side-channel information [3], or were limited to extracting only a small number of layers [15]. A turning point came with the work of Carlini et al. [8], who showed strong connections between the extraction of ReLU-based neural networks and the cryptanalysis of block ciphers. More precisely, they exhibited in this work a differential-style attack capable of recovering significantly larger models than prior approaches. This attack can be divided into two parts: signature extraction (recovering the absolute value of the internal weights) and sign extraction (recovering the signs of the internal weights). This method permitted the authors of [8] to successfully reverse-engineer model parameters trained on random data with at most three hidden layers.
On the other hand, the sign recovery process of [8] was identified to be very time-consuming (exponential in the number of neurons) and therefore impractical for deeper models trained on standard datasets. To overcome this limitation, Canales-Martinez et al. [6] proposed a polynomial-time sign-extraction algorithm, the neuron wiggle.
Recently, Foerster et al. [13] sped up the sign recovery and conducted a thorough performance evaluation of the signature [8] and sign [6] extraction methods. They showed that the signature extraction algorithm does not give satisfying results for deeper layers, even if ran for a much longer time.
In parallel with these efforts to improve performance under full access to confidence scores, other members of the research community have turned their attention to the more realistic “hard-label” scenario [10, 7], where the DNN outputs only the predicted class label, corresponding to the highest confidence score, while the actual confidence values remain hidden. Finally, we emphasize that neither [8, 13] nor [10, 7] managed to extract weights deeper than the third layer (see previous footnote).
We refer to [22] for a recent survey of the field.
@@ 3 Preliminaries
In this section, we introduce important definitions and notations (Section 3.1), followed by assumptions regarding the attack setting (Section 3.2 and Section 3.3) and finally an overview of the original attack of [8] (Section 3.4).
@@ 3.1 Notations and Definitions
An $r$-deep fully connected neural network of architecture $[d_{0},\cdots,d_{r+1}]$ is a function $f:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{r+1}}$ composed of alternating linear layers $\ell^{(i)}:\mathbb{R}^{d_{i-1}}\to\mathbb{R}^{d_{i}}$ and a non-linear activation function $\sigma$ acting component-wise such that: $f=\ell^{(r+1)}\circ\sigma\circ\dots\circ\sigma\circ\ell^{(2)}\circ\sigma\circ\ell^{(1)}$, where $\ell^{(i)}({\bm{x}})={\bm{A}}^{(i)}(x)+b^{(i)}$; ${\bm{A}}^{(i)}\in\mathbb{R}^{{*}^{d_{i}\times d_{i-1}}}$, $b^{(i)}\in\mathbb{R}^{d_{i}}$. We call:
@@
  • ${\bm{A}}^{(i)}$: the $i$-th linear layer weight matrix.
@@
To extend the notations, the architecture $[d_{0},\dots,d_{r+1}]$ can also be written as $d_{0}-\dots-d_{r+1}$. For consecutive layers with the same dimension, an exponential notation can be used for compactness, e.g., $120-10^{(3)}-$ 1 represents the architecture $20-10-10-10-1$.
As in [8, 6, 13], this research only considers fully connected neural networks using the widespread ${\rm ReLU}$ activation function  [21] applied component-wise. An example of such a network is depicted in Figure 1. The structure of a fully connected neural network resembles that of substitution-permutation networks (SPNs) such as the AES [2]. The component-wise application of the activation function is analogous to the use of S-boxes in SPNs, hence the analogy with cryptanalysis introduced in [8].
@@
$\sigma(v)$ $=(\text{ReLU}(v_{1}),\text{ReLU}(v_{2}),\dots,\text{ReLU}(v_{n}))$
@@
$x$$\ell^{(1)}$$\sigma^{(1)}$$\ell^{(2)}$$\sigma^{(2)}$$\ell^{(r)}$$\sigma^{(r)}$$\ell^{(r+1)}$$f(x)$
Figure 1: Fully connected ReLU neural network
In the following, we define a reduced-round neural network $F^{(i)}$ as a function $F^{(i)}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{i+1}}$ given by:
$F^{(i)}=\ell^{(i)}\circ\sigma\circ\dots\circ\sigma\circ\ell^{(2)}\circ\sigma\circ\ell^{(1)}.$
$F^{(i)}$ shares the same linear transformations and activation functions as the original function $f$, up to layer $i$. It will be used to track the transformation of an input vector as it goes through $f$. Observe that $F^{(r+1)}=f$.
More formally, the $k$-th neuron of layer $i$ is defined as the function $\eta^{(i)}_{k}(x)=A^{(i)}_{k}\cdot x+b^{(i)}_{k}$, where $A^{(i)}_{k}$ is the $k$-th row of ${\bm{A}}^{(i)}$ and $b^{(i)}_{k}$ is the $k$-th coordinate of $b^{(i)}$.
A central notion in the analysis of [8] is that of critical points—input points at which a specific neuron becomes inactive (i.e. when its output is zero). These points play a crucial role, as they enable the extraction of the corresponding neuron’s weights through carefully chosen queries. A formal definition is provided below.
@@
An input point $x\in\mathbb{R}^{d_{0}}$ is called a critical point if there exists an index pair $(i,k)$ such that:
$[F^{(i)}(x)]_{k}=0,\quad\text{for }i\in\{1,\ldots,r\},\quad k\in\{1,\ldots,d_{i}\}.$
where $[\cdot]_{k}$ refers to the $k$-th entry of a vector. Equivalently, this holds if $\eta^{(i)}_{k}\circ\sigma\circ F^{(i-1)}(x)=0.$ For this reason, we say that $x$ is a critical point of neuron $\eta^{(i)}_{k}$.
The activation pattern for a given input $x$ is the set of all (active) neurons that contribute to the output $f(x)$. A formal definition is given below:
@@
The activation pattern of $x$ through $f$ is the set
$\mathcal{S}_{f}(x):=\{\eta_{k}^{(i)}\>\big{|}\>[F^{(i)}(x)]_{k}\geq 0,\quad 1\leq i\leq r,1\leq k\leq d_{i}\}.$
If $\eta_{k}^{(i)}\in S_{f}(x)$ then we say that neuron $\eta_{k}^{(i)}$ is active, otherwise we say that it is inactive (for $x$ and $f$).
@@
The polytope of $x$ is the set:
$\mathcal{P}_{x}:=\{x^{\prime}\in\mathbb{R}^{d_{0}}\>|\>\mathcal{S}_{f}(x)=\mathcal{S}_{f}(x^{\prime})\}$
Therefore, $\mathcal{P}_{x}$ represents the region of the input space consisting of all points $x^{\prime}$ close enough to $x$ for the network to activate the same subset of neurons for $x^{\prime}$ as it does for $x$. Later, we will justify the use of the term polytope to describe this space.
@@
First, for $x\in\mathbb{R}^{d_{0}}$, the network $f$ is affine on $\mathcal{P}_{x}$, meaning that there exists $(\Gamma_{x},\gamma_{x})$ such that $\forall x^{\prime}\in\mathcal{P}_{x}$, we have $f(x^{\prime})=\Gamma_{x}x^{\prime}+\gamma_{x}.$ We note:
  • $f_{x}$ the function $x^{\prime}\mapsto\Gamma_{x}x^{\prime}+\gamma_{x}$.
@@
Though ReLUs are not linear, they are piecewise linear. This means that when moving near a non-critical point, all ReLUs remain linear. Recall that critical points are solutions of an equation of the form $A^{(i)}_{k}\cdot F^{(i-1)}(x)+b^{(i)}_{k}=0$. Thus, they form hyperplanes of dimensions $d_{0}-1$, which partition the input space into these different affine regions, hence referred to as polytopes.
We denote by $I_{x}^{(i)}$ the diagonal matrix representing which neurons in layer $i$ are active for the input $x$: the $k$-th coefficient of the diagonal is $1$ if the $k$-th neuron in layer $i$ is active, and 00 otherwise. Thus,
@@
$=I_{x}^{(i)}A^{(i)}\cdots I_{x}^{(2)}A^{(2)}I_{x}^{(1)}A^{(1)}x+\gamma^{(i)}_{x}=\Gamma^{(i)}_{x}x+\gamma^{(i)}_{x}$

### v2

@@ Abstract
Neural network model extraction has recently emerged as an important security concern, as adversaries attempt to recover a network’s parameters via black-box queries. Carlini et al. proposed in CRYPTO’20 a model extraction approach inspired by differential cryptanalysis, consisting of two steps: signature extraction, which extracts the absolute values of network weights layer by layer, and sign extraction, which determines the signs of these signatures. However, in practice this signature-extraction method is limited to very shallow networks only, and the proposed sign-extraction method is exponential in time. Recently, Canales-Martínez et al. (Eurocrypt’24) proposed a polynomial-time sign-extraction method, but it assumes the corresponding signatures have already been successfully extracted and can fail on so-called low-confidence neurons.
In this work, we first revisit and refine the signature extraction process by systematically identifying and addressing for the first time critical limitations of Carlini et al.’s signature-extraction method. These limitations include rank deficiency and noise propagation from deeper layers. To overcome these challenges, we propose efficient algorithmic solutions for each of the identified issues, greatly improving the capabilities of signature extraction. Our approach permits the extraction of much deeper networks than previously possible.
In addition, we propose new methods to improve numerical precision in signature extraction, and enhance the sign extraction part by combining two polynomial methods to avoid exponential exhaustive search in the case of low-confidence neurons. This leads to the very first end-to-end model extraction method that runs in polynomial time.
We validate our attack through extensive experiments on ReLU-based neural networks, demonstrating significant improvements in extraction depth. For instance, our attack extracts consistently at least eight layers of neural networks trained on either the MNIST or CIFAR-10 datasets, while previous works could barely extract the first three layers of networks of similar width. Our results represent a crucial step toward practical attacks on larger and more complex neural network architectures.
@@ 1 Introduction
Neural Networks (NNs) are a class of machine learning models composed of layers of interconnected units (or neurons) that transform input data and learn to recognize patterns through a training process. Deep Neural Networks (DNNs) are a subset of NNs with multiple hidden layers, enabling them to model highly complex functions. DNNs have become indispensable in various applications, including computer vision (e.g., image recognition, and video analysis), natural language processing, automated medical diagnostics, and fraud detection, among others. However, training DNNs requires large datasets, significant computational resources, and carefully fine-tuned algorithms [5, 26]. As a result, trained DNNs have become valuable intellectual assets, making them attractive targets for attackers seeking to extract the model rather than invest in training their own.
@@
In many cases, DNNs are deployed as cloud-based or online services, allowing users to interact with them without direct access to their internal parameters [1, 24]. A natural question, therefore, is how well different DNNs resist attacks in this black-box setting, where an adversary can query the network using random or carefully chosen inputs, potentially adaptively, and exploit the observed outputs to reconstruct either the exact internal parameters or an approximation sufficient to construct a functionally equivalent network.
Neural network model extraction is an old and well-studied problem. The earliest work dates back to 1994, when Fefferman [11] proved that the output of a sigmoid network uniquely determines its architecture and weights, up to trivial equivalences. Later, in 2005, Lowd and Meek [17] introduced new algorithms for reverse engineering linear classifiers and applied these techniques to spam filtering, marking one of the first practical extraction attacks. The field saw renewed interest in 2016 with the work of Tramèr et al. [25], who demonstrated attacks on deployed machine learning models accessible through APIs that output high-precision confidence values (also referred to as scores), in addition to class labels. By exploiting these confidence scores, they successfully mounted attacks on various model types, including logistic regression and decision trees.
For non-linear models, [25] and later [22] introduced the notion of task accuracy extraction, where the goal is to build a model that performs well on the same decision task, achieving high accuracy on predictions, without necessarily reproducing the exact outputs of the original model. This is different from functionally equivalent extraction, where the objective is to replicate the original model’s outputs on all inputs, regardless of the ground truth. Jagielski et al. [14], who introduced this taxonomy, argued that learning-based approaches like [22, 25] are fundamentally limited when it comes to achieving functionally equivalent extraction. Early attempts for functionally equivalent extraction relied on access to internal gradients of the model [19], side-channel information [3], or were limited to extracting only a small number of layers [14].
@@
A significant step in model extraction came in 2020 with Carlini et al. [8], who studied ReLU networks in the S5 setting of [7], which exposes the network’s pre-normalization outputs. In this work, the authors reframed the problem from a cryptanalytic perspective by observing that DNNs share key similarities with block ciphers, including their iterative structure, alternating linear and non-linear layers, and the use in parallel of a small function (S-box in block ciphers, activation function in neural networks) to ensure non-linearity. Taking inspiration from cryptanalytic attacks on block ciphers, particularly differential cryptanalysis [4], Carlini et al. proposed a new two-step iterative approach for recovering a model’s internal parameters. Their method decomposes the extraction process into two distinct phases: signature extraction and sign extraction. In the first step, the absolute values of the weights in a layer are reconstructed (signature extraction), followed by a second step where the correct signs are determined (sign extraction). This process is applied iteratively, proceeding layer by layer. This approach enabled them to successfully extract neural networks, including a model with 100,000100{,}000 parameters trained on the MNIST digit recognition task, using only $2^{21.5}$ queries and less than an hour of computation. However, their method for extracting the signs had an exponential complexity in the number of neurons, limiting its feasibility to small networks with typically very few (at most 3) layers.
@@
In 2024, Canales-Martínez et al. [6] extended the work of Carlini et al. [8] by introducing several new algorithms for the sign-extraction step in ReLU-based neural networks. Their improved sign-extraction methods permitted them to extract the parameters of significantly larger networks than those considered in [8]. For instance, they reported successfully extracting the parameters of a neural network with 8 hidden layers of 256 neurons each (amounting to 1.2 million parameters) trained on the CIFAR-10 dataset [15] for image classification across 10 categories. It is important to note, however, that their experiments assumed the absolute values of the weights had already been recovered using the method of Carlini et al. [8] and were not implemented as full end-to-end attacks.
@@
More recently, Foerster et al. [12] conducted a detailed analysis of the signature and sign extraction procedures, implementing and benchmarking them on various ReLU-based neural networks. Their work revealed that, contrary to what was previously believed, the main practical bottleneck preventing these attacks from advancing beyond the early layers lies in the signature-extraction step introduced in [8]. Because of this, both [8] and [12] failed in practice to recover weight parameters beyond the third layer. Although the authors of [6] reported successful extraction from deeper networks, this was assuming that signatures were already extracted with [8]’s method, which could not work on deeper parts of the networks. In addition, Foerster et al. [12] also identified a critical limitation in the sign-extraction method of [6]: for many networks, the confidence in the recovered signs of neurons does not increase even with additional iterations, contradicting the claim in [6] that wrong signs could be fixed by testing more critical points. To address this, they proposed applying exhaustive search, similar to [8], for neurons with low confidence. This indicates that, for many networks, sign extraction has not yet reached true polynomial-time extraction, leaving it as another unresolved bottleneck.
In parallel with these efforts to improve performance under full access to confidence scores, other members of the research community have turned their attention to the more realistic “hard-label” scenario [7, 9, 29], (S1 setting in the taxonomy of [7]) where the DNN outputs only the predicted class label, corresponding to the highest confidence score, while the actual confidence values remain hidden. We show however that important problems persist even in the more informative S5 setting, and resolving them is a prerequisite for full extraction of deep networks. Moreover, because many techniques are shared across settings, advances achieved in S5 naturally carry over and improve extraction in the more restrictive scenarios as well.
@@
To date, no method can efficiently recover a network’s parameters beyond the third hidden layer in any setting (from S1 to S5) [^1] The authors of [7] report having successfully extracted the 4th layer of a single network. However, in that network, this 4th layer is very contractive, causing the issues identified in our work not to arise.. This highlights the need for significant improvements in the extraction to enable attacks on deeper networks, as well as an exploration of more challenging end-to-end attacks (attacks recovering a model’s parameters solely from black-box input–output pairs, without any simplification such as artificially separating sign and signature extraction).
@@
Prior attacks do not scale to deep neural networks not merely because increased depth raises computational cost, but because their methods are incomplete and overlook issues that grow more severe in deeper models. In this work, we diagnose these challenges and propose practical solutions, thereby enabling effective end-to-end extraction on much deeper neural networks.
Reaching deeper layers. By identifying and addressing two critical limitations in signature extraction which arise as network depth increases, we allow signature extraction to progress beyond the first few layers:
  • Incorrect signature extraction due to rank-deficient systems. The original attack produces incorrect signatures at deeper layers when the linear system used to compute signatures is rank-deficient. We provide a method to increase the rank, yielding correct signature extraction on deeper networks.
  • Overlooked influence of deeper layers. Contrary to what was previously believed, we show that signatures originating from deeper layers than the one attacked can be misinterpreted as coming from the target layer. We propose strategies to filter them out depending on the neural network targeted.
End-to-end extraction. We present the first end-to-end model extraction that runs in polynomial time, achieved through notable enhancements to signature precision improvement and sign-extraction methods.
  • Signature precision improvement. In addition to Carlini et al. [8]’s approach for improving signature precision, we identify and address two additional sources of imprecision in signature extraction, which improves extraction efficiency and supports wide layers.
  • Confident sign extraction. As noted by [12], the neuron wiggle method of [6] fails for the networks we consider on so-called low-confidence neurons. Instead of the exponential exhaustive search that [12] uses, we combine two polynomial methods from [6] to avoid relying on low-confidence neurons.
We validate all our improvements through extensive experiments on ReLU-based neural networks, showing significant improvements in extraction depth. Notably, we present the first end-to-end model extraction on eight-layer networks with eight neurons per layer trained on MNIST and CIFAR-10 datasets, extracting almost all the weights and achieving a very low relative error $\epsilon\leq 3.6\times 10^{-4}$ on $73\%$ of the input space, while previous works could barely extract the first three layers. These results mark a significant step toward practical attacks on larger, more complex neural network architectures. Our full code is accessible in https://github.com/PsyduckLiu/End-to-End-Deep-Neural-Network-Extraction.
The rest of the article is organized as follows. Section 2 introduces preliminaries: notations, the attack model, and a short overview of the original signature-extraction and sign-extraction methods. Section 3 presents our improvements to signature extraction. Section 4 describes our new sign-extraction strategy. Section 5 introduces our evaluation metrics and discusses a taxonomy of unrecovered weights. Section 6 reports our experimental evaluation and end-to-end results. We evaluate our attack on three different networks: one trained on CIFAR-10 with architecture $3072\text{-}8^{(8)}\text{-}1$ and two networks trained on MNIST with architectures $784\text{-}8^{(8)}\text{-}1$ and $784\text{-}16^{(8)}\text{-}1$, which we refer to as Model I, II, and III, respectively. Finally, Section 7 concludes and discusses limitations.
@@ 2 Preliminaries
In this section, we introduce important definitions and notations (Section 2.1), followed by assumptions regarding the attack setting and goal (Section 2.2) and finally an overview of the original attack of [8] and [6] (Section 2.3).
@@ 2.1 Notations and Definitions
An $r$-deep fully connected neural network of architecture $d_{0},\cdots,d_{r+1}]$ is a function $f:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{r+1}}$ composed of alternating linear layers $\ell^{(i)}:\mathbb{R}^{d_{i-1}}\to\mathbb{R}^{d_{i}}$ and a non-linear activation function $\sigma$ acting component-wise such that: $f=\ell^{(r+1)}\circ\sigma\circ\dots\circ\sigma\circ\ell^{(2)}\circ\sigma\circ\ell^{(1)}$, where $\ell^{(i)}(x)=A^{(i)}(x)+b^{(i)}$; $A^{(i)}\in\mathbb{R}^{{d_{i}\times d_{i-1}}}$, $b^{(i)}\in\mathbb{R}^{d_{i}}$. We call:
@@
  • $A^{(i)}$: the $i$-th linear layer weight matrix.
@@
To extend the notations, the architecture $d_{0},\dots,d_{r+1}]$ can also be written as $d_{0}-\dots-d_{r+1}$. For consecutive layers with the same dimension, an exponential notation can be used for compactness, e.g., $20-10^{(3)}-1$ represents the architecture 20101010120-10-10-10-1.
As in [6, 8, 12], this research only considers fully connected neural networks using the widespread $\rm ReLU$} activation function [20] applied component-wise. The structure of a fully connected neural network resembles that of substitution-permutation networks (SPNs) such as the AES [2]. The component-wise application of the activation function is analogous to the use of S-boxes in SPNs, hence the analogy with cryptanalysis introduced in [8].
@@
$\sigma(v)$ $=({\rm ReLU}(v_{1}),{\rm ReLU}(v_{2}),\dots,{\rm ReLU}(v_{n}))$
@@
In the following, we define a reduced-round neural network $F^{(i)}$ for $1\leq i\leq r$ as a function $F^{(i)}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{i}}$ given by:
$F^{(i)}=\sigma\circ\ell^{(i)}\circ\sigma\circ\dots\circ\sigma\circ\ell^{(2)}\circ\sigma\circ\ell^{(1)}.$
$F^{(i)}$ shares the same linear transformations and activation functions as the original function $f$, up to layer $i$. It will be used to track the transformation of an input vector as it goes through $f$. By $F^{(r+1)}$ and $F^{(0)}$, we mean respectively $f$ and the identity function.
The $k$-th neuron of layer $i$ is defined as the function $\eta^{(i)}_{k}(x)=A^{(i)}_{k}\cdot x+b^{(i)}_{k}$, where $A^{(i)}_{k}$ is the $k$-th row of $A^{(i)}$ and $b^{(i)}_{k}$ is the $k$-th coordinate of $b^{(i)}$.
A central notion in the analysis of [8] is that of critical points of neurons. They enable the extraction of their corresponding neuron’s weights through carefully chosen queries.
@@
An input point $x\in\mathbb{R}^{d_{0}}$ is called a critical point of a neuron $\eta_{k}^{(i)}$ if
$\eta_{k}^{(i)}\circ F^{(i-1)}(x)=0$
The activation pattern for a given input $x$ is the set of all (active) neurons that contribute to the output $f(x)$.
@@
The activation pattern of $x$ through $f$ at layer $i$, for $1\leq i\leq r$, is the set
$\mathcal{S}^{(i)}_{f}(x):=\{\eta_{k}^{(j)}\>\big|\>\eta_{k}^{(j)}\circ F^{(j-1)}(x)>0,\quad 1\leq j\leq i,1\leq k\leq d_{j}\}.$
If $i=r$, we note $\mathcal{S}^{(i)}_{f}$ as $\mathcal{S}_{f}$ and refer to it as the activation pattern of $x$ through $f$. If $\eta_{k}^{(i)}\in\mathcal{S}_{f}(x)$ then we say that neuron $\eta_{k}^{(i)}$ is active, otherwise we say that it is inactive (for $x$ and $f$).
@@
The polytope of $x$ at layer $i$, $\mathcal{P}^{(i)}_{x}$, is the largest connected open subset of
$\{x^{\prime}\in\mathbb{R}^{d_{0}}\>|\>\mathcal{S}^{(i)}_{f}(x)=\mathcal{S}^{(i)}_{f}(x^{\prime})\}$
containing $x$. If $i=r$, we note $\mathcal{P}^{(i)}_{x}$ as $\mathcal{P}_{x}$ and refer to it as the polytope of $x$.
Therefore, $\mathcal{P}_{x}$ represents the region of the input space consisting of all points $x^{\prime}$ connected to $x$ that have the same activation pattern. Let’s now justify the use of the term polytope to describe this space.
@@
First, for $x\in\mathbb{R}^{d_{0}}$, for all $1\leq i\leq r$, the network $F^{(i)}$ is affine on $\mathcal{P}^{(i)}_{x}$, meaning that there exists $\Gamma_{x}^{(i)}\in\mathbb{R}^{d_{i}\times d_{0}}$, $\gamma^{(i)}_{x}\in\mathbb{R}^{d_{i}}$ such that $\forall x^{\prime}\in\mathcal{P}^{(i)}_{x}$, we have $F^{(i)}(x^{\prime})=\Gamma^{(i)}_{x}x^{\prime}+\gamma^{(i)}_{x}$. We note:
@@
Though ReLUs are not linear, they are piecewise linear. This means that when moving near a non-critical point, all ReLUs act as linear functions. Recall that critical points are solutions of an equation of the form $A^{(i)}_{k}\cdot F^{(i-1)}(x)+b^{(i)}_{k}=0$. Thus, they form hyperplanes which partition the input space into these different affine regions, hence referred to as polytopes.
We denote by $I_{x}^{(i)}$ the diagonal matrix representing which neurons in layer $i$ are active for the input $x$: the $k$-th coefficient of the diagonal is $1$ if the $k$-th neuron in layer $i$ is active, and 0 otherwise. Thus,
@@
$=I_{x}^{(i)}A^{(i)}\cdots I_{x}^{(2)}A^{(2)}I_{x}^{(1)}A^{(1)}x+\gamma^{(i)}_{x}=\Gamma^{(i)}_{x}x+\gamma^{(i)}_{x}$.